# Origin and emergent features of many-body dynamical localization

### v2

@@ Abstract
The question of whether interactions can break dynamical localization in quantum kicked rotor systems has been the subject of a long-standing debate. Here, we introduce an extended mapping from the kicked Lieb-Liniger model to a high-dimensional lattice model and reveal universal features: on-site pseudorandomness and hybrid exponential–algebraic decay couplings with increasing momenta. We find that the exponent and the amplitude of the algebraic decay undergo a crossover as the interaction strength increases. This mapping predicts the existence of dynamical localization and its breakdown at large kick strengths and intermediate interaction strengths. An analysis of the generalized fractal dimension and level-spacing ratio supports these findings, highlighting the presence of near integrability and multifractality in different regions of parameter space. Our results offer an explanation for the occurrence of many-body dynamical localization, particularly in strongly correlated quantum gases, and are anticipated to generalize to systems of many particles.
@@
Introduction—Quantum coherence can give rise to captivating phenomena. For example, the quantum kicked rotor (QKR), as a driven single-particle system, exhibits dynamical localization (DL) [1, 2, 3, 4, 5, 6]: non-resonant periodic driving paradoxically leads to suppressed energy absorption. Such an unexpected halt in energy growth in the quantum realm starkly contrasts our everyday experience, which tells us that driven systems generally thermalize to infinite temperature. DL for the QKR can be understood as Anderson localization (AL) in momentum space, characterized by a freezing of the momentum distribution in the course of its evolution [7, 8]. This behavior has been studied extensively in cold-atom experiments over the past three decades [9, 10, 11, 12, 13, 14, 15, 16]. However, in reality, particles interact with each other. This raises a fundamental question: Can driven quantum systems avoid thermalization in the presence of inter-particle interactions? This challenge has spurred the development and investigation of various many-body QKR models, where particle-particle correlations are introduced through mechanisms such as periodic kicks [17, 18, 19, 20] or static contact interactions [21, 22, 23]. These models predict a diverse range of behavior, including both DL and delocalized phases. Understanding the mechanism of localization sheds light on the role of many-body interactions in driven quantum systems.
@@
The QKR model with contact interactions has received particular attention since it can be realized in cold-atom platforms. Predictions based on the mean-field approximation suggest that the system delocalizes at long times with a sub-diffusive behavior of the energy [24, 25, 26, 27], consistent with recent experimental observations [28, 29]. However, in one dimension, the mean-field approach runs into severe limitations [30, 31]. Beyond this approximation, numerical studies treating two interacting bosons have yielded contradictory results [23, 32]. In the Tonks-Girardeau (TG) regime, non-perturbative techniques such as the Bose-Fermi mapping [33, 34, 35, 36, 37, 38] predict that DL persists in the interacting many-body situation. This phenomenon has been called many-body dynamical localization (MBDL) [21, 22, 39]. It has been observed in recent experimental work [40]. The momentum distribution was shown to freeze over the entire range from weak to strong interactions despite the strong drive.
@@
To date, most studies on MBDL have focused on obtaining the time evolution of the energy and the momentum distribution, starting from low-energy initial states [21, 22, 39]. However, the microscopic mechanism driving MBDL remains little understood. In this paper, we investigate MBDL in the Lieb-Liniger (LL) model [41, 42, 43, 44, 45] augmented by a kicking term [21]. Using the LL eigenstates as a basis, we establish an extended mapping from this quantum system to a $N$-dimensional lattice model with hybrid exponential–algebraic decay couplings, in which the algebraic couplings undergo a crossover as the interaction strength increases. Our analysis predicts localization and its breakdown. Our conclusion is supported by inspecting the generalized fractal dimensions (GFD) and the energy-level-spacing statistics.
@@
The parameters $g$ and $K$ denote the interaction strength and the kick strength, respectively. For convenience, we represent time in units of $T$, position in units of $1/k_{\mathrm{L}}$, and momentum in units of $\hbar_{\rm{eff}}=4T\hbar k_{\mathrm{L}}^{2}/m_{\mathrm{A}}$, which is an effective Planck constant [46]. The various limits are well understood: For K=0K\!=\!0, we recover the integrable LL model, which can be solved exactly via the Bethe-Ansatz technique [41, 42]. The LL eigenstates $\lvert\mathcal{I}\rangle=\lvert m_{1}m_{2}\cdots m_{N}\rangle$ are labeled by $N$ quasi-momenta $\{m_{j}\}$ with the eigenenergy being $E_{\mathcal{I}}=(\hbar_{\mathrm{eff}}^{2}/2)\sum_{j=1}^{N}m_{j}^{2}$. The LL parameter $\gamma=gL/(N\hbar^{2}_{\rm{eff}})$ represents the dimensionless interaction strength. For $K\neq 0$, it is known that DL and MBDL occur in the two limits g=0g\!=\!0 and $g\!=\!\infty$ [6, 21]. In both cases the problem can be mapped onto a single-particle problem. In the intermediate interaction regime, it is unclear whether the localization phenomenon persists, as illustrated in Fig 1(b).
@@
and $\Pi_{\mathcal{I}}$ denotes the coefficient containing the information about the Floquet eigenstate. The diagonal elements $V_{\mathcal{I}}$ are pseudorandom since $E_{\mathcal{I}}$ has a quadratic dependence on quasi-momentum. The off-diagonal elements $W_{\mathcal{I}\mathcal{I}^{\prime}}$ represent the couplings between Bethe eigenstates in a $N$-dimensional quasi-momentum space spanned by $\{m_{j}|j=1,2,\dots,N\}$.
Figure 2: (a-c) The off-diagonal term $\lvert W_{0\mathcal{I}}\rvert$ as a function of energy $E_{\mathcal{I}}-E_{0}$ for (a) $g=10^{-1}$, (b) $g=10^{1}$, and (c) $g=10^{3}$ in log-log scale (blue dots). The black lines stand for the Bethe-Ansatz results. The black circles and squares represent the off-diagonal term for the mapped free bosons and fermions, respectively. The green and red dashed lines denote algebraic decay scaling as ${E_{\mathcal{I}}}^{-2}$ and ${E_{\mathcal{I}}}^{-3/2}$, respectively. The insets show the histograms of the on-site term $V_{\mathcal{I}}$, fitted by a Lorentzian distribution. The momentum cut-off, the kick strength, and the particle number are $M=100$, $K=0.5$ and $N=2$, respectively. (d-f) show the same data as in (a-c) but for $N=3$ and $M=35$. (g, h) show the decay exponent $\lambda$ and amplitude $A$ as a function of $g$, respectively.
We firstly employ ED to obtain the LL eigenstates for $N=2,3$. Starting from the ground state of the LL model, we compute the off-diagonal term $W_{0\mathcal{I}}$ as a function of the energy difference $(E_{\mathcal{I}}-E_{0})$ between an excited state $|\mathcal{I}\rangle$ and the ground state $|0\rangle$ for varying interaction strengths, as shown in Fig. 2(a-f). Notably, for both $N=2$ and $3$, the off-diagonal terms consistently exhibit hybrid decay: exponential-like decay at low energies and algebraic decay at high energies. Compared to the $N=2$ case, more branches of algebraic decay emerge for $N=3$. We notice that this exponential-like decay is the one that one would expect for free bosons at $g\rightarrow 0$ or for fermions at $g=\infty$. The pseudorandom on-site terms consistently follow a Lorentzian distribution (see insets). Remarkably, the decay exponent and amplitude of the algebraic tail have a nontrivial dependence on the interaction strength. We further fit the top branch of these algebraic tails of $\lvert W_{0\mathcal{I}}\rvert$ to a power-law function $AE_{\mathcal{I}}^{-\lambda/2}$, and we report the exponent $\lambda$ and the amplitude $A$ for different $g$ in Fig. 2 (g) and (h). As the interaction strength increases, the decay exponent $\lambda$ shows a crossover from $\lambda=4$ to a lower value around $3$. In parallel, the amplitude $A$ exhibits a non-monotonic change. This indicates that the variation of the algebraic tails is driven by interactions.
@@
$W_{\mathcal{II^{\prime}}}\sim\left\{\begin{aligned} &\frac{\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)}{4\pi^{2}A_{g}}\cdot\frac{1}{q^{4}}&&\mathrm{for}\kern 5.0pt2qA_{g}\rightarrow\infty,\,q\gg q^{\prime},\\ &\frac{\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)q^{\prime}A_{g}}{\pi^{2}}\cdot\frac{1}{q^{3}}&&\mathrm{for}\kern 5.0pt2qA_{g}\rightarrow 0,\,\ \ q\gg q^{\prime},\end{aligned}\right$. (6)
where $\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)\!=\!\frac{K}{2\hbar_{\mathrm{eff}}}\int_{0}^{2\pi}\frac{\cos{(x_{\mathrm{c}})}\exp\{2i(Q-Q^{\prime})x_{\mathrm{c}}\}}{\cos^{2}{[K\cos{(x_{\mathrm{c}})}/\hbar_{\mathrm{eff}}]}}dx_{\mathrm{c}}$, $A_{g}\!=\!\hbar^{2}_{\mathrm{eff}}/g$, and $x_{\mathrm{c}}=(x_{1}+x_{2})/2$ is the center-of-mass (CM) position. Q(q)Q\ (q) and $Q^{\prime}\ (q^{\prime})$ are the CM (relative) momentum of eigenstates $\mathcal{I}$ and $\mathcal{I^{\prime}}$, respectively. The kicks nontrivially couple to the hopping in the CM momentum $\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)$, which decays exponentially with $|Q-Q^{\prime}|$ and indicates DL in the CM momentum space. Generally, this is expected for any $N$ since $\hat{H}_{\mathrm{I}}$ conserves the CM momentum [46]. In the relative momentum, there is indeed a crossover in the tail $q^{-\lambda}$ of the hopping $W_{\mathcal{II^{\prime}}}$ from $\lambda=4$ to $\lambda=3$ and the tail’s amplitude is first proportional, then inversely proportional to $g$ as $g$ grows. As shown in Fig. 2, the Bethe-Ansatz results are in agreement with the ED results. Importantly, Eq. (6) reveals the mechanism: the kicks lead to exponential decay between LL eigenstates from different CM momentum sectors, while interactions further generate an algebraic tail in the relative momentum.
@@
where $c^{s,PP^{\prime}}_{h,j}=\sum_{l=j}^{N}\Delta m^{PP^{\prime}}_{l}\!\!+\!s\cdot H(h-j)$ with $\Delta m^{PP^{\prime}}_{l}\!=\!m_{P(l)}\!-\!m^{\prime}_{P^{\prime}(l)}$ and $H(\cdot)$ being the Heaviside step function. The Bethe coefficient is $A_{P}\propto(-1)^{\mathrm{sgn}(P)}\prod_{j>l}[m_{P(j)}-m_{P(l)}-ig]$ with $P$ denoting a permutation operation of $N$ quasi-momenta $\{m_{j}\}$. Here, the kicks effectively couple two different Bethe states whose CM momenta satisfy $\sum_{l=1}^{N}\Delta m^{PP^{\prime}}_{l}=2\pi/L$, leading to a 2nd-order pole $c^{s,PP^{\prime}}_{h,1}=c^{s,PP^{\prime}}_{h,N+1}$. For stronger kicks, one only needs to account for additional CM momentum sectors regardless of $N$. Moreover, consideration of more particles only introduces additional first-order poles, since $\Delta m^{PP^{\prime}}_{l}\neq 0$ generally holds for different Bethe states. This analysis suggests that the algebraic tail is universal for arbitrary $N$.
@@
Figure 3: The decay exponent $\lambda$ (a) and the amplitude AA (b) as a function of $\gamma$ for different NN. The inset of (a) illustrates the typical two-particle excitations $\{m_{1},m_{N}\}$ of the excited state $\lvert\mathcal{I}\rangle$, while the other quasi-momenta $\{m_{2}\cdots m_{N-1}\}$ are kept identical to their ground-state value $\lvert 0\rangle$.
@@
For increasing values of $N$, we compute the off-diagonal terms $W_{0\mathcal{I}}$ using Eq. (7), and then fit $W_{0\mathcal{I}}$ to an algebraic function $A{\delta m}^{-\lambda}$ [46] with $\delta m$ representing the standard deviation of the quasi-momentum of the excited state $\lvert\mathcal{I}\rangle$. For a typical algebraic tail (see the top branch in Fig. 2), only two momenta $m_{1}$ and $m_{N}$ vary under the condition $m_{1}+m_{N}=2\pi/L$ and $\sum^{N-1}_{j=2}m_{j}=0$, as illustrated in the inset of Fig. 3(a). We plot the decay exponent and the amplitude as a function of $\gamma$ in Fig. 3. The decay exponents for different $N$ converge, exhibiting the identical transition from $4$ to $3$ as seen in Eq. (6). The amplitude $A$ changes non-monotonically with increasing $\gamma$.
Generally, for $\lvert\mathcal{I}\rangle$ in each CM momentum sector (fixed $\sum^{N}_{j=1}m_{j}$), the algebraically decaying branches of $W_{0\mathcal{I}}$ are characteristic of a two-particle excitation (e.g., varying $\{m_{1},m_{N}\}$ with a different set of fixed quasi-momenta $\{m_{2},\dots,m_{N-1}\}$) [46]. Also, the off-diagonal term exhibits an anisotropy between excited Bethe states $|\mathcal{I}\rangle$ and $|\mathcal{I}^{\prime}\rangle$ [46].
Localization and its breakdown—In the basis of Bethe states, the mapped model forms a complex quantum network characterized by a set of degrees of freedom $\{m_{1},\dots,m_{N}\}$. Its diagonal elements are pseudorandom, while the off-diagonal elements exhibit exponential-algebraic decay along each $m_{j}$. Driven by a finite interaction $g$, particularly at an intermediate strength, the algebraic tails emerge in the relative momenta, whereas the diagonal pseudorandomness is only weakly sensitive to variations of the interaction strength. Furthermore, as the kick strength increases, the amplitude of these algebraic tails grows and the long-range couplings become more pronounced. This undermines the localization effect. Consequently, a breakdown of localization occurs at large kick strengths and intermediate interaction strengths, corresponding to the maximal amplitude of the algebraically decaying tails. Note that, if the kick strength is sufficiently small, localization persists for all interaction strengths for finite particle numbers.
In the two limits $g=0$ and $\infty$, the state $\lvert\mathcal{I}\rangle$ reduces to the free boson wavefunction and the Slater’s determinant with a symmetrizer, respectively. In both limits, the algebraically decaying tails vanish, making the model equivalent to $N$ independent copies of 1D Anderson-like models [2, 8] in momentum space. Localization is expected to occur since $V_{{m}}$ are pseudorandom numbers drawn from a Lorentzian distribution and the off-diagonal term decays sufficiently fast [49]. Thus, in these two limits, localization is expected for arbitrary kick strengths.
@@
Statistics of eigenstates—Here, we apply the traditional diagnostic to the Floquet eigenspectrum and eigenstates to verify the above predictions for the mapped model. We diagonalize $\hat{U}$, focus on the mid spectrum $(\theta_{\alpha}=0)$, typically extracting $500$ eigenstates for various values of the cutoff $M$, ranging from $10$ to $32$. We first study the Floquet eigenstates based on the participation entropies (PE) $S_{\beta}$. Given an eigenstate $\lvert\phi\rangle$ and a $\mathcal{N}$-dimensional basis $\{\lvert\alpha\rangle\}$, the PE are defined as [50, 51, 52, 53] $S_{\beta}=\mathrm{ln}\left(\sum_{\alpha=1}^{\mathcal{N}}\lvert\phi_{\alpha}\rvert^{2\beta}\right)/(1-\beta)$ with $\lvert\phi\rangle=\sum_{\alpha=1}^{\mathcal{N}}\phi_{\alpha}\lvert\alpha\rangle$. Then one can define the GFD $D_{\beta}$ as $D_{\beta}=\lim_{\mathcal{N}\rightarrow\infty}S_{\beta}/\mathrm{ln}\mathcal{N}$. If $D_{\beta}$ nontrivially depends on $\beta$, the state is multifractal, whereas for constant $D_{\beta}$, the state is fractal [51]. For a fully delocalized state, $D_{\beta}=1$, while $D_{\beta}=0$ represents the scenario of Anderson localizion. Values of $0<D_{\beta}<1$, indicate an extended but nonergodic state.

### v3

@@ Abstract
The question of whether interactions can break dynamical localization in quantum kicked rotor systems has been the subject of a long-standing debate. Here, we introduce an extended mapping from the kicked Lieb-Liniger model to a high-dimensional lattice model and reveal universal features: on-site pseudorandomness and hybrid exponential–algebraic decay couplings with increasing momenta. We find that the exponent and the amplitude of the algebraic decay undergo a crossover as the interaction strength increases. This mapping uncovers the origin of dynamical localization and the interaction effect on the integrability of the system. An analysis of the generalized fractal dimension and level-spacing ratio supports these findings, highlighting the presence of near integrability and multifractality in different regions of parameter space. Our results offer an explanation for the occurrence of many-body dynamical localization, particularly in strongly correlated quantum gases, and are anticipated to generalize to systems of many particles.
@@
Introduction—Quantum coherence can give rise to captivating phenomena. For example, the quantum kicked rotor (QKR), as a driven single-particle system, exhibits dynamical localization (DL) [1, 2, 3, 4, 5, 6]: nonresonant periodic driving paradoxically leads to suppressed energy absorption. Such an unexpected halt in energy growth in the quantum realm starkly contrasts our everyday experience, which tells us that driven systems generally thermalize to infinite temperature. DL for the QKR can be understood as Anderson localization (AL) in momentum space, characterized by a freezing of the momentum distribution in the course of its evolution [7, 8]. This behavior has been studied extensively in cold-atom experiments over the past three decades [9, 10, 11, 12, 13, 14, 15, 16]. However, in reality, particles interact with each other. This raises a fundamental question: Can driven quantum systems avoid thermalization in the presence of interparticle interactions? This challenge has spurred the development and investigation of various many-body QKR models, where particle-particle correlations are introduced through mechanisms such as periodic kicks [17, 18, 19, 20] or static contact interactions [21, 22, 23]. These models predict a diverse range of behavior, including both DL and delocalized phases. Understanding the mechanism of localization sheds light on the role of many-body interactions in driven quantum systems.
@@
The QKR model with contact interactions has received particular attention since it can be realized in cold-atom platforms. Predictions based on the mean-field approximation suggest that the system delocalizes at long times with a subdiffusive behavior of the energy [24, 25, 26, 27], consistent with recent experimental observations [28, 29]. However, in one dimension, the mean-field approach runs into severe limitations [30, 31]. Beyond this approximation, numerical studies treating two interacting bosons have yielded contradictory results [23, 32]. In the Tonks-Girardeau (TG) regime, nonperturbative techniques such as the Bose-Fermi mapping [33, 34, 35, 36, 37, 38] predict that DL persists in the interacting many-body situation. This phenomenon has been called many-body dynamical localization (MBDL) [21, 22, 39]. It has been observed in recent experimental work [40]. The momentum distribution was shown to freeze over the entire range from weak to strong interactions despite the strong drive.
@@
To date, most studies on MBDL have focused on obtaining the time evolution of the energy and the momentum distribution, starting from low-energy initial states [21, 22, 39]. However, the microscopic mechanism driving MBDL remains little understood. In this Letter, we investigate MBDL in the Lieb-Liniger (LL) model [41, 42, 43, 44, 45] augmented by a kicking term [21]. Using the LL eigenstates as a basis, we establish an extended mapping from this quantum system to a NN-dimensional lattice model with hybrid exponential–algebraic decay couplings, in which the algebraic couplings undergo a crossover as the interaction strength increases. Our analysis reveals the origin of localization and the key factors underlying its potential breakdown. Our conclusion is supported by inspecting the generalized fractal dimensions (GFD) and the energy-level-spacing statistics.
@@
The parameters $g$ and $K$ denote the interaction strength and the kick strength, respectively. For convenience, we represent time in units of $T$, position in units of $1/k_{\mathrm{L}}$, and momentum in units of $\hbar_{\rm{eff}}=4T\hbar k_{\mathrm{L}}^{2}/m_{\mathrm{A}}$, which is an effective Planck constant [46]. The various limits are well understood: for K=0K\!=\!0, we recover the integrable LL model, which can be solved exactly via the Bethe-Ansatz technique [41, 42]. The LL eigenstates $\lvert\mathcal{I}\rangle=\lvert m_{1}m_{2}\cdots m_{N}\rangle$ are labeled by $N$ quasimomenta $\{m_{j}\}$ with the eigenenergy being $E_{\mathcal{I}}=(\hbar_{\mathrm{eff}}^{2}/2)\sum_{j=1}^{N}m_{j}^{2}$. The LL parameter $\gamma=gL/(N\hbar^{2}_{\rm{eff}})$ represents the dimensionless interaction strength. For $K\neq 0$, it is known that DL and MBDL occur in the two limits g=0g\!=\!0 and $g\!=\!\infty$ [6, 21]. In both cases the problem can be mapped onto a single-particle problem. In the intermediate interaction regime, it is unclear whether the localization phenomenon persists, as illustrated in Fig 1(b).
@@
and $\Pi_{\mathcal{I}}$ denotes the coefficient containing the information about the Floquet eigenstate. The diagonal elements $V_{\mathcal{I}}$ are pseudorandom since $E_{\mathcal{I}}$ has a quadratic dependence on quasimomentum. The off-diagonal elements $W_{\mathcal{I}\mathcal{I}^{\prime}}$ represent the couplings between Bethe eigenstates in an $N$-dimensional quasimomentum space spanned by $\{m_{j}|j=1,2,\dots,N\}$.
Figure 2: (a-c) The off-diagonal term $\lvert W_{0\mathcal{I}}\rvert$ as a function of energy $E_{\mathcal{I}}-E_{0}$ for (a) $g=10^{-1}$, (b) $g=10^{1}$, and (c) $g=10^{3}$ in log-log scale (blue dots). The black lines stand for the Bethe-Ansatz results. The black circles and squares represent the off-diagonal term for the mapped free bosons and fermions, respectively. The green and red dashed lines denote algebraic decay scaling as ${E_{\mathcal{I}}}^{-2}$ and ${E_{\mathcal{I}}}^{-3/2}$, respectively. The insets show the histograms of the on-site term $V_{\mathcal{I}}$, fitted by a Lorentzian distribution. The momentum cutoff, the kick strength, and the particle number are $M=100$, $K=0.5$ and $N=2$, respectively. (d-f) show the same data as in (a-c) but for $N=3$ and $M=35$. (g, h) show the decay exponent $\lambda$ and amplitude $A$ as a function of $g$, respectively.
We first employ ED to obtain the LL eigenstates for $N=2,3$. Starting from the ground state of the LL model, we compute the off-diagonal term $W_{0\mathcal{I}}$ as a function of the energy difference $(E_{\mathcal{I}}-E_{0})$ between an excited state $|\mathcal{I}\rangle$ and the ground state $|0\rangle$ for varying interaction strengths, as shown in Fig. 2(a-f). Notably, for both $N=2$ and $3$, the off-diagonal terms consistently exhibit hybrid decay: exponential-like decay at low energies and algebraic decay at high energies. Compared to the $N=2$ case, more branches of algebraic decay emerge for $N=3$. We notice that this exponential-like decay is the one that one would expect for free bosons at $g\rightarrow 0$ or for fermions at $g=\infty$. The pseudorandom on-site terms consistently follow a Lorentzian distribution (see insets). Remarkably, the decay exponent and amplitude of the algebraic tail have a nontrivial dependence on the interaction strength. We further fit the top branch of these algebraic tails of $\lvert W_{0\mathcal{I}}\rvert$ to a power-law function $AE_{\mathcal{I}}^{-\lambda/2}$, and we report the exponent $\lambda$ and the amplitude $A$ for different $g$ in Fig. 2 (g) and (h). As the interaction strength increases, the decay exponent $\lambda$ shows a crossover from $\lambda=4$ to a lower value around $3$. In parallel, the amplitude $A$ exhibits a nonmonotonic change. This indicates that the variation of the algebraic tails is driven by interactions.
@@
$W_{\mathcal{II^{\prime}}}\sim\left\{\begin{aligned} &\frac{\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)}{4\pi^{2}A_{g}}\frac{1}{q^{4}}&&\mathrm{for}\kern 5.0pt2qA_{g}\rightarrow\infty,\,q\gg q^{\prime},\\ &\frac{\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)q^{\prime}A_{g}}{\pi^{2}}\frac{1}{q^{3}}&&\mathrm{for}\kern 5.0pt2qA_{g}\rightarrow 0,\,\ \ q\gg q^{\prime},\end{aligned}\right$. (6)
where $\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)\!=\!\frac{K}{2\hbar_{\mathrm{eff}}}\int_{0}^{2\pi}\frac{\cos{(x_{\mathrm{c}})}\exp\{2i(Q-Q^{\prime})x_{\mathrm{c}}\}}{\cos^{2}{[K\cos{(x_{\mathrm{c}})}/\hbar_{\mathrm{eff}}]}}dx_{\mathrm{c}}$, $A_{g}\!=\!\hbar^{2}_{\mathrm{eff}}/g$, and $x_{\mathrm{c}}=(x_{1}+x_{2})/2$ is the center-of-mass (c.m.) position. Q(q)Q\ (q) and $Q^{\prime}\ (q^{\prime})$ are the c.m. (relative) momentum of eigenstates $\mathcal{I}$ and $\mathcal{I^{\prime}}$, respectively. The kicks nontrivially couple to the hopping in the c.m. momentum $\mathcal{B}_{Q-Q^{\prime}}^{(2)}(0)$, which decays exponentially with $|Q-Q^{\prime}|$ and indicates DL in the c.m. momentum space. Generally, this is expected for any $N$ since $\hat{H}_{\mathrm{I}}$ conserves the c.m. momentum [46]. In the relative momentum, there is indeed a crossover in the tail $q^{-\lambda}$ of the hopping $W_{\mathcal{II^{\prime}}}$ from $\lambda=4$ to $\lambda=3$ and the tail’s amplitude is first proportional, then inversely proportional to $g$ as $g$ grows. As shown in Fig. 2, the Bethe-Ansatz results are in agreement with the ED results. Importantly, Eq. (6) reveals the mechanism: the kicks lead to exponential decay between LL eigenstates from different c.m. momentum sectors, while interactions further generate an algebraic tail in the relative momentum.
@@
where $c^{s,PP^{\prime}}_{h,j}=\sum_{l=j}^{N}\Delta m^{PP^{\prime}}_{l}\!\!+\!s\cdot H(h-j)$ with $\Delta m^{PP^{\prime}}_{l}\!=\!m_{P(l)}\!-\!m^{\prime}_{P^{\prime}(l)}$ and $H(\cdot)$ being the Heaviside step function. The Bethe coefficient is $A_{P}\propto(-1)^{\mathrm{sgn}(P)}\prod_{j>l}[m_{P(j)}-m_{P(l)}-ig]$ with $P$ denoting a permutation operation of $N$ quasimomenta $\{m_{j}\}$. Here, the kicks effectively couple two different Bethe states whose c.m. momenta satisfy $\sum_{l=1}^{N}\Delta m^{PP^{\prime}}_{l}=2\pi/L$, leading to a second-order pole $c^{s,PP^{\prime}}_{h,1}=c^{s,PP^{\prime}}_{h,N+1}$. For stronger kicks, one only needs to account for additional c.m. momentum sectors regardless of $N$. Moreover, consideration of more particles only introduces additional first-order poles, since $\Delta m^{PP^{\prime}}_{l}\neq 0$ generally holds for different Bethe states. This analysis suggests that the algebraic tail is universal for arbitrary $N$.
@@
Figure 3: The decay exponent $\lambda$ (a) and the amplitude AA (b) as a function of $\gamma$ for different NN. The inset of (a) illustrates the typical two-particle excitations $\{m_{1},m_{N}\}$ of the excited state $\lvert\mathcal{I}\rangle$, while the other quasimomenta $\{m_{2}\cdots m_{N-1}\}$ are kept identical to their ground-state value $\lvert 0\rangle$. (c) IPR of eigenstates of the mapped model as a function of $\gamma$ for different $K$, which has been averaged over $500$ eigenstates around midspectrum. The momentum cutoff and the particle number are respectively $M=35$ and $N=3$.
@@
For increasing values of $N$, we compute the off-diagonal terms $W_{0\mathcal{I}}$ using Eq. (7), and then fit $W_{0\mathcal{I}}$ to an algebraic function $A{\delta m}^{-\lambda}$ [46] with $\delta m$ representing the standard deviation of the quasimomentum of the excited state $\lvert\mathcal{I}\rangle$. For a typical algebraic tail (see the top branch in Fig. 2), only two momenta $m_{1}$ and $m_{N}$ vary under the condition $m_{1}+m_{N}=2\pi/L$ and $\sum^{N-1}_{j=2}m_{j}=0$, as illustrated in the inset of Fig. 3(a). We plot the decay exponent and the amplitude as a function of $\gamma$ in Fig. 3. The decay exponents for different $N$ converge, exhibiting the identical transition from $4$ to $3$ as seen in Eq. (6). The amplitude $A$ changes nonmonotonically with increasing $\gamma$. Generally, for $\lvert\mathcal{I}\rangle$ in each c.m. momentum sector (fixed $\sum^{N}_{j=1}m_{j}$), the algebraically decaying branches of $W_{0\mathcal{I}}$ are characteristic of a two-particle excitation (e.g., varying $\{m_{1},m_{N}\}$ with a different set of fixed quasimomenta $\{m_{2},\dots,m_{N-1}\}$) [46]. Also, the off-diagonal term exhibits an anisotropy between excited Bethe states $|\mathcal{I}\rangle$ and $|\mathcal{I}^{\prime}\rangle$ [46]. To further demonstrate the effect of the algebraic tail on localization, we show the inverse participation ratio (IPR) $\sum_{\mathcal{I}}\lvert\Pi_{\mathcal{I}}\rvert^{4}$ of the eigenstates of Eq. (4) in Fig. 3(c). All the data consistently exhibit a significant dip (gray shaded) around intermediate $\gamma$, indicating the extended property of eigenstates. Larger kicks greatly decrease the IPR value. Therefore, the mapped model will first transition toward ergodicity at large kicks and intermediate interaction strength.
Localization and its breakdown— The mapping approach uncovers crucial insights about localization. The mapped model forms a complex $N$-dimensional quantum network characterized by a set of degrees of freedom $\{m_{1},\dots,m_{N}\}$. Its diagonal elements exhibit a consistently pseudorandom distribution, constituting the fundamental origin of localization in analogy to the Anderson problem. Interaction induces nontrivial algebraic decay into the off-diagonal elements along specific directions in the high-dimensional network, this famously can thwart localization. Particularly at intermediate interaction strengths, the off-diagonal elements attain maximum amplitude and exhibit the slowest decay, as shown in Fig. 3(c). Within this region, large kicks will eventually render the system ergodic. Eq. (7) can be equally expressed in terms of the form factor $W_{\mathcal{I}\mathcal{I}^{\prime}}\propto\langle\mathcal{I}\lvert\Psi^{\dagger}(x)\Psi(x)\rvert\mathcal{I}^{\prime}\rangle$, where $\Psi(x)$ is the field operator [46]. A well-known exact form of the matrix element is available [49, 50, 51, 52]. Recent work shows that, in the thermodynamic limit, matrix element of the density operator $\Psi^{\dagger}(0)\Psi(0)$ between two Bethe states with extensive excitations decays superexponentially in the system size [53]. This makes the high-dimensional network effectively disconnected in most directions. We therefore expect that, in the thermodynamic limit, localization will survive, at small $K$ and any value of $g$.
@@
Statistics of eigenstates—Here, we apply the traditional diagnostic to the Floquet eigenspectrum and eigenstates to verify the above predictions for the mapped model. We diagonalize $\hat{U}$, focus on the mid spectrum $(\theta_{\alpha}=0)$, typically extracting $500$ eigenstates for various values of the cutoff $M$, ranging from $10$ to $32$. We first study the Floquet eigenstates based on the participation entropies (PE) $S_{\beta}$. Given an eigenstate $\lvert\phi\rangle$ and an $\mathcal{N}$-dimensional basis $\{\lvert\alpha\rangle\}$, the PE are defined as [54, 55, 56, 57] $S_{\beta}=\mathrm{ln}\left(\sum_{\alpha=1}^{\mathcal{N}}\lvert\phi_{\alpha}\rvert^{2\beta}\right)/(1-\beta)$ with $\lvert\phi\rangle=\sum_{\alpha=1}^{\mathcal{N}}\phi_{\alpha}\lvert\alpha\rangle$. Then one can define the GFD $D_{\beta}$ as $D_{\beta}=\lim_{\mathcal{N}\rightarrow\infty}S_{\beta}/\mathrm{ln}\mathcal{N}$. If $D_{\beta}$ nontrivially depends on $\beta$, the state is multifractal, whereas for constant $D_{\beta}$, the state is fractal [55]. For a fully delocalized state, $D_{\beta}=1$, while $D_{\beta}=0$ represents the scenario of Anderson localizion. Values of $0<D_{\beta}<1$, indicate an extended but nonergodic state.